# Comparing Bayesian and Frequentist Inference in Biological Models: A Comparative Analysis of Accuracy, Uncertainty, and Identifiability

### v1

@@ 3.0.0.1 LV model
where $\alpha$ is the intrinsic growth rate of the prey population, $\beta$ is the predation rate coefficient, $\gamma$ is the mortality rate of the predator population, and $\delta$ is the predator’s growth rate per prey consumed. The initial conditions $x_{0}$ and $y_{0}$ specify the starting population sizes of prey and predator, respectively. The LV model serves as a classical benchmark in population ecology, offering a well-understood test case for examining parameter identifiability and uncertainty propagation under full versus partial observability.
@@
Figure 2: LV predator–prey diagram. Circles represent the prey $x$ and predator $y$ populations. The self-loop on $x$ indicates intrinsic growth at rate $\alpha$. Curved arrows between $x$ and $y$ represent interactions: prey loss due to predation and predator growth from consuming prey. The downward arrow on $y$ represents natural mortality.
@@ 3.0.0.2 Generalized Logistic Model
where $C(t)$ is the cumulative number of cases at time $t$, $r$ is the generalized growth rate, $K$ is the final epidemic size, and $p\in[0,1]$ controls the growth dynamics: $p=0$ corresponds to constant incidence, $0<p<1$ to sub-exponential growth, and $p=1$ to exponential growth. Here, $C_{0}$ represents the first reported case. Moreover, $\frac{dC}{dt}$ serves as the observation operator mapping the model states to the reported incidence data. The GLM was chosen because it provides a parsimonious, single-equation description of epidemic trajectories and serves as a tractable case for testing inference accuracy and uncertainty when the full time series is directly observed. This model has also been widely applied in epidemiological settings to capture early and full-phase epidemic dynamics chowell2016zika . The GLM has been widely used as a phenomenological growth model to capture epidemic curves with rapid increase and saturation chowell2016zika; pell2018ebola .
@@ 4 Methods
This section describes the Bayesian and Frequentist estimation frameworks used for parameter inference, uncertainty quantification, and forecasting. Let $Y=(y_{t_{1}},\ldots,y_{t_{n}})$ denote the observed data at discrete time points $t_{1},\ldots,t_{n}$, and let $\boldsymbol{\theta}$ represent the vector of unknown model parameters to be estimated. For the Lotka–Volterra (LV) model, $\boldsymbol{\theta}=(\alpha,\beta,\gamma,\delta)$. For the generalized logistic model (GLM), $\boldsymbol{\theta}=(r,p,K)$. For the SEIUR model, $\boldsymbol{\theta}=(\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1})$. Both estimation approaches assume a normal (Gaussian) error structure to ensure comparability of results.
@@ 4.1 Bayesian Inference
Bayesian inference integrates prior knowledge with observed data to obtain the posterior distribution of model parameters vanDeSchoot2021. This probabilistic framework is particularly effective when prior information is available or when data are sparse, noisy, or subject to partial observability, as it provides comprehensive uncertainty quantification through posterior distributions gelman2020bayesian; grinsztajn2021bayesian.
@@ 4.1.2 Likelihood Function
We assume that observation errors are independent and identically distributed (i.i.d.) following a normal distribution with constant variance. For the LV model with observed prey and predator populations, the likelihood is:
@@
independently for $j=1,\ldots,n$, where $x_{\boldsymbol{\theta}}(t)$ and $y_{\boldsymbol{\theta}}(t)$ denote the prey and predator population solutions to the LV ODE system given parameters $\boldsymbol{\theta}$, and $\sigma_{\text{prey}}^{2}$ and $\sigma_{\text{predator}}^{2}$ are observation error variances.
@@ 4.1.3 Prior Distributions
Prior distributions encode existing knowledge or beliefs about parameter values before observing data greenland2009bias . For parameters with established biological or epidemiological interpretations, we specify informative priors based on previous studies. When prior knowledge is limited, we use weakly informative or uniform priors over plausible parameter ranges. For structural identifiability analysis with simulated data, priors were centered on true parameter values with varying degrees of precision to assess the impact of prior strength on parameter recovery. Table 14,18,19,23,24,28,29 summarizes the parameter bounds used for each model. The normal error variance $\sigma^{2}$ is assigned a weakly informative inverse-gamma or half-Cauchy prior to allow the data to inform the observation error scale.
@@ 4.1.4 Posterior Sampling and Convergence Diagnostics
Posterior distributions are approximated using Markov Chain Monte Carlo (MCMC) sampling implemented through the Hamiltonian Monte Carlo (HMC) algorithm in Stan vehtari2021rhat; burkner2017brms. For each model, we ran four independent chains with a sufficient number of iterations (typically 2,000–4,000) after a warm-up (burn-in) period to ensure convergence to the stationary distribution. Convergence was assessed using the Gelman–Rubin diagnostic $\hat{R}$ statistic burkner2017brms, which compares between-chain and within-chain variance; values of $\hat{R}<1.05$ (ideally $\hat{R}\approx 1.01$) indicate successful convergence. Effective sample size (ESS) was monitored to ensure adequate sampling of the posterior distribution. Trace plots and posterior density plots were visually inspected to confirm mixing and convergence.
Posterior summaries are reported as medians with 95% credible intervals (CrIs), which represent the central 95% of the posterior distribution. For forecasting, we propagate parameter uncertainty by sampling parameters values from the posterior distribution, solving the ODE for each sample, and summarizing the resulting forecast distribution with median and 95% prediction intervals (PIs).
@@ 4.1.5 Computational Implementation
Bayesian inference was conducted using the (BFF) toolbox Karami2025BayesianFitForecast, an R package designed for fitting and forecasting ODE-based epidemic models. The BFF toolbox provides an automated workflow that generates Stan code based on user-specified model structures, priors, and data inputs, eliminating the need for users to program directly in Stan. The toolbox outputs include posterior parameter distributions, convergence diagnostics ($\hat{R}$, ESS), trace plots, posterior density plots, and forecasting results with quantified uncertainty. Performance metrics including mean absolute error (MAE), mean squared error (MSE), weighted interval score (WIS), and 95% PI coverage are automatically computed to evaluate model fit and forecast accuracy.
@@ 4.2 Frequentist Inference
Frequentist estimation treats model parameters as fixed but unknown quantities and estimates them by optimizing an objective function based solely on the observed data, without incorporating prior information chowell2017primer; chowell2020ensembles . Uncertainty quantification is performed through bootstrap resampling procedures.
@@ 4.2.1 Parameter Estimation
where $\mu_{j}(\boldsymbol{\theta})$ is the model-predicted value at time $t_{j}$ given parameters $\boldsymbol{\theta}$. For the LV model, $\mu_{j}(\boldsymbol{\theta})$ consists of predicted prey and predator populations $(x_{\boldsymbol{\theta}}(t_{j}),y_{\boldsymbol{\theta}}(t_{j}))$. For the GLM, $\mu_{j}(\boldsymbol{\theta})=dC_{\boldsymbol{\theta}}(t_{j})/dt$. For the SEIUR model, $\mu_{j}(\boldsymbol{\theta})=\kappa\rho E_{\boldsymbol{\theta}}(t_{j})$.
The optimization is performed using gradient-based algorithms such as the Levenberg–Marquardt algorithm or trust-region methods, which iteratively search the parameter space to minimize the objective function. The observed data variance is estimated as:
@@ 4.2.2 Uncertainty Quantification via Parametric Bootstrap
To quantify parameter uncertainty and construct confidence intervals, we employ a parametric bootstrap procedure chowell2017primer; banks2014uncertainty. This approach generates synthetic datasets by resampling from the fitted model and re-estimates parameters for each synthetic dataset. The distribution of bootstrap parameter estimates characterizes sampling variability. The steps are as follows:
4.2.2.1 Generate bootstrap samples:
For each bootstrap replicate $b=1,\ldots,B$ (typically B=200B=200500500), generate a synthetic dataset $\{y_{t_{1}}^{b},\ldots,y_{t_{n}}^{b}\}$ by sampling from the fitted model:
@@ 4.2.2.1 Generate bootstrap samples:
4.2.2.2 Re-estimate parameters:
For each bootstrap sample $\{y_{t_{1}}^{b},\ldots,y_{t_{n}}^{b}\}$, solve the NLS optimization problem (Eq. 11) to obtain bootstrap parameter estimates $\hat{\boldsymbol{\theta}}^{b}$.
4.2.2.3 Construct confidence intervals:
Use the empirical distribution of $\{\hat{\boldsymbol{\theta}}^{b}:b=1,\ldots,B\}$ to construct confidence intervals. For example, the 95% confidence interval for parameter $\theta_{k}$ is given by the 2.5th and 97.5th percentiles of $\{\hat{\theta}_{k}^{b}:b=1,\ldots,B\}$.
@@ 4.2.3 Computational Implementation
Frequentist inference was implemented using the QuantDiffForecast (QDF) MATLAB toolbox Chowell2024QuantDiffForecast, which provides a comprehensive framework for parameter estimation, uncertainty quantification, and forecasting for ODE models. The QDF toolbox supports multiple optimization algorithms, flexible error structures (normal, Poisson, negative binomial), user-defined ODE systems, and automated bootstrap-based uncertainty quantification. The toolbox outputs parameter estimates with confidence intervals, fitted model trajectories, forecast distributions, and performance metrics (MAE, MSE, WIS, 95% PI coverage).
@@ 4.3.3 Coverage of the 95% Prediction Interval
where $L_{t_{i}}$ and $U_{t_{i}}$ are the lower and upper bounds of the 95% prediction interval at time $t_{i}$, and $\mathbf{1}(\cdot)$ is an indicator function equal to 1 if the condition is true and 0 otherwise. Ideally, the coverage should be close to 95%. Coverage significantly below 95% indicates overly narrow prediction intervals (underestimation of uncertainty), while coverage above 95% suggests overly wide intervals (overestimation of uncertainty).
@@ 4.3.4 Weighted Interval Score (WIS)
WIS is a proper scoring rule that evaluates the quality of the entire predictive distribution by combining sharpness (interval width) and calibration (penalties for observations outside intervals) (Gneiting2007; bracher2021evaluating). The interval score (IS) for a single prediction interval at level $\alpha$ is:
@@
4.3.4.1 Sharpness
(ul)(u-l), the width of the central $(1-\alpha)\times 100\%$ prediction interval. Narrower intervals receive lower scores, rewarding precise forecasts.
4.3.4.2 Underprediction penalty
$\frac{2}{\alpha}(l-y)\mathbf{1}(y<l)$, penalizes observations falling below the lower bound $l$, with the penalty proportional to the distance (ly)(l-y).
4.3.4.3 Overprediction penalty
$\frac{2}{\alpha}(y-u)\mathbf{1}(y>u)$, penalizes observations exceeding the upper bound $u$, with the penalty proportional to the distance (yu)(y-u).
To comprehensively evaluate the full predictive distribution, we compute the WIS by averaging interval scores over multiple prediction interval levels $(1-\alpha_{1})<(1-\alpha_{2})<\cdots<(1-\alpha_{K})$ along with the predictive median $\tilde{y}$ (which can be viewed as a prediction interval at level $(1-\alpha_{0})\to 0$):
@@ 5 Structural Identifiability
Structural identifiability (SI) analysis determines whether model parameters can, in principle, be uniquely recovered from perfect, noise-free observations of the system outputs, given the model structure and observation scheme (raue2009structural; identifiability_review; stigter2023structural). A parameter is structurally identifiable if its value can be uniquely determined from the input-output relationship defined by the model equations and the set of observable variables. Conversely, a parameter is structurally unidentifiable if multiple distinct parameter values produce identical model outputs, making unique parameter recovery impossible regardless of data quality or quantity.
Structural identifiability is a prerequisite for meaningful parameter estimation: if a parameter is structurally unidentifiable, no amount of data or sophisticated inference methods can recover its true value (practical_identifiability; chowell2023structural). SI analysis is particularly important in compartmental models where only a subset of state variables is observed, as partial observability often leads to identifiability deficits (dankwa2022identifiability). Understanding which parameters are identifiable under different observation scenarios guides experimental design, informs prior specification in Bayesian inference, and helps interpret estimation results.
@@ 5.1 Methodology
We conducted symbolic structural identifiability analysis using StructuralIdentifiability.jl (stigter2023structural), a Julia package that employs differential algebra methods to determine identifiability of parameters in ODE models. For each model, we analyzed two scenarios: (1) known initial conditions (ICs), where the initial values of all state variables are assumed to be known exactly, and (2) unknown initial conditions, where ICs are treated as additional unknown parameters to be estimated. The distinction between known and unknown ICs is critical, as identifiability results can differ substantially depending on whether ICs must be inferred from data (identifiability_methods).
For the Lotka–Volterra (LV) model, we examined three observation schemes to assess how partial observability affects identifiability:
5.1.0.1 LV-1 (Both prey and predator observed):
Both $x(t)$ (prey) and $y(t)$ (predator) populations are observed.
5.1.0.2 LV-2 (Predator only):
Only $y(t)$ (predator) is observed; prey population $x(t)$ is unobserved.
5.1.0.3 LV-3 (Prey only):
Only $x(t)$ (prey) is observed; predator population $y(t)$ is unobserved.

### v2

@@ 3.0.0.1 Lotka-Volterra model
where $\alpha$ is the intrinsic growth rate of the prey population, $\beta$ is the predation rate coefficient, $\gamma$ is the mortality rate of the predator population, and $\delta$ is the predator’s growth rate per prey consumed. The initial conditions $x_{0}$ and $y_{0}$ specify the starting population sizes of prey and predator, respectively. The Lotka–Volterra model serves as a classical benchmark in population ecology, offering a well-understood test case for examining parameter identifiability and uncertainty propagation under full versus partial observability.
@@
Figure 2: Lotka–Volterra predator–prey diagram. circles represent the prey $x$ and predator $y$ populations. The self-loop on $x$ indicates intrinsic growth at rate $\alpha$. Curved arrows between $x$ and $y$ represent interactions: prey loss due to predation and predator growth from consuming prey. The downward arrow on $y$ represents natural mortality.
@@ 3.0.0.2 Generalized Logistic Model
where $C(t)$ is the cumulative number of cases at time $t$, $r$ is the generalized growth rate, $K$ is the final epidemic size, and $p\in[0,1]$ controls the growth dynamics: $p=0$ corresponds to constant incidence, $0<p<1$ to sub-exponential growth, and $p=1$ to exponential growth. Here, $C_{0}$ represents the first reported case. Moreover, $\frac{dC}{dt}$ serves as the observation operator mapping the model states to the reported incidence data. The GLM was chosen because it provides a parsimonious, single-equation description of epidemic trajectories and serves as a tractable case for testing inference accuracy and uncertainty when the full time series is directly observed. This model has also been widely applied in epidemiological settings to capture early and full-phase epidemic dynamics [85] . The GLM has been widely used as a phenomenological growth model to capture epidemic curves with rapid increase and saturation [85, 86] .
@@ 3.0.0.3 SEIUR Model
The time-dependent transmission rate $\beta_{f}(t)$ is modeled in a piecewise manner to reflect changes in population behavior and control measures following an intervention at time $t_{\mathrm{int}}$. Before the intervention, transmission is assumed to be constant at level $\beta_{0}$, representing relatively uncontrolled epidemic spread. After $t_{\mathrm{int}}$, transmission decreases smoothly toward a lower asymptotic level $\beta_{1}$ through an exponential decay, capturing the gradual adoption and increasing effectiveness of interventions such as mobility reduction, masking, or behavioral adaptation. The decay rate $q_{1}$ controls how rapidly these effects take hold, allowing the model to represent both abrupt and slow transitions while remaining parsimonious and identifiable.
@@ 4 Methods
This section describes the Bayesian and Frequentist estimation frameworks used for parameter inference, uncertainty quantification, and forecasting. Let $Y=(y_{t_{1}},\ldots,y_{t_{n}})$ denote the observed data at discrete time points $t_{1},\ldots,t_{n}$, and let $\boldsymbol{\theta}$ represent the vector of unknown model parameters to be estimated. For the Lotka–Volterra (LV) model, $\boldsymbol{\theta}=(\alpha,\beta,\gamma,\delta)$. For the generalized logistic model (GLM), $\boldsymbol{\theta}=(r,p,K)$. For the SEIUR model, $\boldsymbol{\theta}=(\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1})$. Both estimation approaches assume a normal error structure to ensure comparability of results.
@@ 4.1 Bayesian Inference
Bayesian inference integrates prior knowledge with observed data to obtain the posterior distribution of model parameters [87]. This probabilistic framework is particularly effective when prior information is available or when data are sparse, noisy, or subject to partial observability, as it provides comprehensive uncertainty quantification through posterior distributions [88, 89].
@@ 4.1.2 Likelihood Function
We assume that observation errors are independent and identically distributed (i.i.d.) following a normal distribution with constant variance. For the Lotka–Volterra model with observed prey and predator populations, the likelihood is:
@@
independently for $j=1,\ldots,n$, where $x_{\boldsymbol{\theta}}(t)$ and $y_{\boldsymbol{\theta}}(t)$ denote the prey and predator population solutions to the Lotka–Volterra ODE system given parameters $\boldsymbol{\theta}$, and $\sigma_{\text{prey}}^{2}$ and $\sigma_{\text{predator}}^{2}$ are observation error variances.
@@ 4.1.3 Prior Distributions
Prior distributions encode existing knowledge or beliefs about parameter values before observing data [44] . For parameters with established biological or epidemiological interpretations, we specify informative priors based on previous studies. When prior knowledge is limited, we use weakly informative or uniform priors over plausible parameter ranges. For structural identifiability analysis with simulated data, priors were centered on true parameter values with varying degrees of precision to assess the impact of prior strength on parameter recovery. Table 28 summarizes the prior distributions used for each model. The normal error variance $\sigma^{2}$ is assigned a weakly informative inverse-gamma or half-Cauchy prior to allow the data to inform the observation error scale.
@@ 4.1.4 Posterior Sampling and Convergence Diagnostics
Posterior distributions are approximated using Markov Chain Monte Carlo (MCMC) sampling implemented through the Hamiltonian Monte Carlo (HMC) algorithm in Stan [62, 65]. For each model, we ran four independent chains with a sufficient number of iterations (typically 2,000–4,000) after a burn-in period to ensure convergence to the stationary distribution. Convergence was assessed using the Gelman–Rubin diagnostic $\hat{R}$ statistic [65], which compares between-chain and within-chain variance; values of $\hat{R}<1.05$ (ideally $\hat{R}\approx 1.01$) indicate successful convergence. Effective sample size (ESS) was monitored to ensure adequate sampling of the posterior distribution. Trace plots and posterior density plots were visually inspected to confirm mixing and convergence.
Posterior summaries are reported as medians with 95% prediction intervals (PIs), which represent the central 95% of the posterior distribution. For forecasting, we propagate parameter uncertainty by sampling parameters values from the posterior distribution, solving the ODE for each sample, and summarizing the resulting forecast distribution with median and 95% prediction intervals (CIs).
@@ 4.1.5 Computational Implementation
Bayesian inference was conducted using the (BFF) toolbox [26], an R package designed for fitting and forecasting ODE-based epidemic models. The BFF toolbox provides an automated workflow that generates Stan code based on user-specified model structures, priors, and data inputs, eliminating the need for users to program directly in Stan. The toolbox outputs include posterior parameter distributions, convergence diagnostics ($\hat{R}$, ESS), trace plots, posterior density plots, and forecasting results with quantified uncertainty. Performance metrics including mean absolute error (MAE), mean squared error (MSE), weighted interval score (WIS), and 95% PI coverage are automatically computed to evaluate model fit and forecast accuracy.
@@ 4.2 Frequentist Inference
Frequentist estimation treats model parameters as fixed but unknown quantities and estimates them by optimizing an objective function based solely on the observed data, without incorporating prior information [33, 34] . Uncertainty quantification is performed through bootstrap resampling procedures.
@@ 4.2.1 Parameter Estimation
where $\mu_{j}(\boldsymbol{\theta})$ is the model-predicted value at time $t_{j}$ given parameters $\boldsymbol{\theta}$. For the Lotka–Volterra model, $\mu_{j}(\boldsymbol{\theta})$ consists of predicted prey and predator populations $(x_{\boldsymbol{\theta}}(t_{j}),y_{\boldsymbol{\theta}}(t_{j}))$. For the GLM, $\mu_{j}(\boldsymbol{\theta})=dC_{\boldsymbol{\theta}}(t_{j})/dt$. For the SEIUR model, $\mu_{j}(\boldsymbol{\theta})=\kappa\rho E_{\boldsymbol{\theta}}(t_{j})$.
In our study, the Frequentist inference is performed using MATLAB’s fmincon with the Sequential Quadratic Programming (SQP) algorithm. To mitigate sensitivity to initial conditions and reduce the risk of convergence to local minima, the optimization is wrapped within a MultiStart framework using a CustomStartPointSet, with an optional RandomStartPointSet.
The observed data variance is estimated as:
@@ 4.2.2 Uncertainty Quantification via Parametric Bootstrap
To quantify parameter uncertainty and construct prediction intervals, we employ a parametric bootstrap procedure [33, 35]. This approach generates synthetic datasets by resampling from the fitted model and re-estimates parameters for each synthetic dataset. The distribution of bootstrap parameter estimates characterizes sampling variability. The steps are as follows:
  1. 1.
    Generate bootstrap samples: For each bootstrap replicate $b=1,\ldots,B$ (typically B=200B=200500500), generate a synthetic dataset $\{y_{t_{1}}^{b},\ldots,y_{t_{n}}^{b}\}$ by sampling from the fitted model:
@@
  2. 2.
    Re-estimate parameters: For each bootstrap sample $\{y_{t_{1}}^{b},\ldots,y_{t_{n}}^{b}\}$, solve the NLS optimization problem (Eq. 11) to obtain bootstrap parameter estimates $\hat{\boldsymbol{\theta}}^{b}$.
  3. 3.
    Construct prediction intervals: Use the empirical distribution of $\{\hat{\boldsymbol{\theta}}^{b}:b=1,\ldots,B\}$ to construct prediction intervals. For example, the 95% prediction interval for parameter $\theta_{k}$ is given by the 2.5th and 97.5th percentiles of $\{\hat{\theta}_{k}^{b}:b=1,\ldots,B\}$.
@@ 4.2.3 Computational Implementation
Frequentist inference was implemented using the QuantDiffForecast (QDF) MATLAB toolbox [25], which provides a comprehensive framework for parameter estimation, uncertainty quantification, and forecasting for ODE models. The QDF toolbox supports multiple optimization algorithms, flexible error structures (normal, Poisson, negative binomial), user-defined ODE systems, and automated bootstrap-based uncertainty quantification. The toolbox outputs parameter estimates with prediction intervals, fitted model trajectories, forecast distributions, and performance metrics (MAE, MSE, WIS, 95% PI coverage).
@@ 4.3.3 Coverage of the 95% prediction interval
where $L_{t_{i}}$ and $U_{t_{i}}$ are the lower and upper bounds of the 95% prediction interval at time $t_{i}$, and $\mathbf{1}(\cdot)$ is an indicator function equal to 1 if the condition is true and 0 otherwise. Ideally, the coverage should be close to 95%. Coverage significantly below 95% indicates overly narrow intervals (underestimation of uncertainty), while coverage above 95% suggests overly wide intervals (overestimation of uncertainty).
@@ 4.3.4 Weighted Interval Score (WIS)
WIS is a proper scoring rule that evaluates the quality of the entire predictive distribution by combining sharpness (interval width) and calibration (penalties for observations outside intervals) [74, 73]. The interval score (IS) for a single prediction interval at level $\alpha$ is:
@@
  • Sharpness: (ul)(u-l), the width of the central $(1-\alpha)\times 100\%$ prediction interval. Narrower intervals receive lower scores, rewarding precise forecasts.
  • Underprediction penalty: $\frac{2}{\alpha}(l-y)\mathbf{1}(y<l)$, penalizes observations falling below the lower bound $l$, with the penalty proportional to the distance (ly)(l-y).
  • Overprediction penalty: $\frac{2}{\alpha}(y-u)\mathbf{1}(y>u)$, penalizes observations exceeding the upper bound $u$, with the penalty proportional to the distance (yu)(y-u).
To comprehensively evaluate the full predictive distribution, we compute the WIS by averaging interval scores over multiple interval levels $(1-\alpha_{1})<(1-\alpha_{2})<\cdots<(1-\alpha_{K})$ along with the predictive median $\tilde{y}$ (which can be viewed as an interval at level $(1-\alpha_{0})\to 0$):
@@ 5 Structural Identifiability
Structural identifiability (SI) analysis determines whether model parameters can, in principle, be uniquely recovered from perfect, noise-free observations of the system outputs, given the model structure and observation scheme [27, 71]. A parameter is structurally identifiable if its value can be uniquely determined from the input-output relationship defined by the model equations and the set of observable variables. Conversely, a parameter is structurally unidentifiable if multiple distinct parameter values produce identical model outputs, making unique parameter recovery impossible regardless of data quality or quantity.
Structural identifiability is a prerequisite for meaningful parameter estimation: if a parameter is structurally unidentifiable, no amount of data or sophisticated inference methods can recover its true value [28, 27] . SI analysis is particularly important in compartmental models where only a subset of state variables is observed, as partial observability often leads to identifiability deficits [72] . Understanding which parameters are identifiable under different observation scenarios guides experimental design, informs prior specification in Bayesian inference, and helps interpret estimation results.
@@ 5.1 Methodology
We conducted symbolic structural identifiability analysis using StructuralIdentifiability.jl [71], a Julia package that employs differential algebra methods to determine identifiability of parameters in ODE models. For each model, we analyzed two scenarios: (1) known initial conditions (ICs), where the initial values of all state variables are assumed to be known exactly, and (2) unknown initial conditions, where ICs are treated as additional unknown parameters to be estimated. The distinction between known and unknown ICs is critical, as identifiability results can differ substantially depending on whether ICs must be inferred from data [70].
For the (LV) model, we examined three observation schemes to assess how partial observability affects identifiability:
  • LV-1 (Both prey and predator observed): Both $x(t)$ (prey) and $y(t)$ (predator) populations are observed.
  • LV-2 (Predator only): Only $y(t)$ (predator) is observed; prey population $x(t)$ is unobserved.
  • LV-3 (Prey only): Only $x(t)$ (prey) is observed; predator population $y(t)$ is unobserved.